OpenStudy (anonymous):

Common factor : (2x+1)3×4(5x−1)3×5+(5x−1)4×3(2x+1)2×2

OpenStudy (anonymous):

\[(2x+1)3×4(5x−1)3×5+(5x−1)4×3(2x+1)2×2\]

myininaya (myininaya):

Is this what you mean? \[(2x+1)^3\cdot 4(5x-1)^3 \cdot 5+(5x-1)^4 \cdot 3(2x+1)^2 \cdot 2 \]

OpenStudy (anonymous):

ohh yes you're right ! it wouldn't let me paste the equation right

myininaya (myininaya):

Ok So we have two terms Do you see what they both have in common?

OpenStudy (anonymous):

so what just \[(2x+1) ^3 \times 4 (5x-1)^3 \times 5 \] has in common ?

myininaya (myininaya):

What does that term you just wrote have in common with that second term ?

OpenStudy (anonymous):

\[(5x-1)^3, (2x+1)^2\] are both in common

myininaya (myininaya):

yes so we can factor that out

myininaya (myininaya):

For example: \[a^3 \cdot 4b^3 \cdot 5+b^4\cdot 3 a^2 \cdot 2\] Both of the terms in this expression have the following factors in common: \[2,a^2,b^3\] So we can factor these common factors out but doing so we just divide each term we factored it from by those factors So like this \[2a^2b^3( \frac{a^3 \cdot 4 b^3 \cdot 5}{2a^2b^3}+\frac{b^4 \cdot 3a^2 \cdot 2}{2a^2b^3})\] \[2a^2b^3(a \cdot 2 \cdot 5+b \cdot 3 )\]

myininaya (myininaya):

And of course you can multiply 2 and 5 to simplify

OpenStudy (anonymous):

here's the actual problem with the solution two on page two. http://sh.triton.net/Math/mcv/Derivatives/chain_rule_short_way.pdf so this would be the chain rule, right ?

OpenStudy (anonymous):

the fact that you changed the variables to a and b

myininaya (myininaya):

I was doing an example

OpenStudy (anonymous):

no i know i was just talking about the method

myininaya (myininaya):

..Do you mean your problem? That is product + chain rule

OpenStudy (anonymous):

okay so for the problem i have so far would this be correct ? \[a^3 \times 4b^3 + b^4 \times 3a^2 \times 2\]

myininaya (myininaya):

Assuming you are using a substitution, no In the first term you are missing a factor 5. Do you need me to check your derivative too?

OpenStudy (anonymous):

oh yeah the five is missing

OpenStudy (anonymous):

so would i need to derive this from here on ?

myininaya (myininaya):

derive?

myininaya (myininaya):

Do you mean factor?

OpenStudy (anonymous):

i'm actually really confused, i dont know what to do from here at all

myininaya (myininaya):

Do you want to factor? Well you told me what those two terms had in common which was 2,a^2,b^3 Correct? Well factor those out like I did above for you

OpenStudy (anonymous):

oh alright 1

OpenStudy (anonymous):

\[a^2 (a+3) + b^3 (b+ 4 \times 5) \]

OpenStudy (anonymous):

is that right ?

myininaya (myininaya):

No I will draw it. Give me a moment.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Omg THANK YOU SO MUCH !

OpenStudy (anonymous):

i have an exam tomorrow and i was rusty on this topic, i think i understand it now

myininaya (myininaya):

Keep practicing :)